# Topology-Informed Graph Transformer

### v2

@@ 1 INTRODUCTION
Transformers have achieved remarkable success in domains such as Natural Language Processing (Vaswani et al., 2023) and Computer Vision (Dosovitskiy et al., 2021). Motivated by their prowess, researchers have applied them to the field of Graph Neural Networks (GNNs). They aimed to surmount the limitations of Message-Passing Neural Networks (MPNNs), a subset of GNNs, by attempting to address challenges such as over-smoothing (Oono & Suzuki, 2021), over-squashing (Alon & Yahav, 2021), and restricted expressive power (Xu et al., 2019; Morris et al., 2021). An exemplary application of the integration of Transformers into GNNs is the Graph Transformer. One way to achieve this integration is by applying multi-head attention mechanism of Transformers to each node of a given graph dataset. Such a technique treats the set of all nodes as being fully interconnected or connected by edges. This approach, however, often has limited capabilities in guaranteeing a strong inductive bias  [^1] For example, unless trained over large scale data, transformers may lack inherent inductive biases in modeling permutation and translational invariance, hierarchical structures, and local data structures  (Hahn, 2020; Dosovitskiy et al., 2021; Xu et al., 2021), hence making it prone to overfitting. To address this drawback, previous studies focused on devising new implementations to blend Graph Transformers with other deep learning techniques including MPNNs, thereby yielding promising empirical results  (Yang et al., 2021; Ying et al., 2021; Dwivedi & Bresson, 2021; Chen et al., 2022; Hussain et al., 2022; Zhang et al., 2023; Ma et al., 2023; Rampášek et al., 2023; Kong et al., 2023; Zhang et al., 2022).
@@
While Graph Transformers boast considerable advancements, enhancing them by strengthening their discriminative powers in distinguishing isomorphism classes of graphs still remains a challenge, an approach that can potentially boost their performance in predicting various properties of graph datasets. For instance, studies based on MPNN have enhanced node attributes by utilizing high-dimensional complexes, persistent homological techniques, and recurring subgraph structures (Carrière et al., 2020; Bodnar et al., 2021b; Bouritsas et al., 2021; Wijesinghe & Wang, 2021; Bevilacqua et al., 2022; Park et al., 2022; Horn et al., 2022; Choi et al., 2023; Li et al., 2023; Choi et al., 2024; Zhou et al., 2024). Other approaches addressing these limitations include incorporating positional encoding grounded in random walk strategies, Laplacian Positional Encoding (PE), node degree centrality, and shortest path distance. Furthermore, structure encoding based on substructure similarity has been introduced to amplify the inductive biases inherent in the Transformer.
@@
This paper introduces a Topology-Informed Graph Transformer (TIGT), which embeds topological inductive biases to augment the model’s expressive power and predictive efficacy. Prior to the Transformer layer, TIGT augments each node attribute with a topological positional embedding layer based on the differences of universal covers (or unfolding trees) obtained from the original graph and collections of unions of its cyclic subgraphs. These new topological biases allow TIGT to encapsulate the first homological invariants (or cyclic structures) [^2] Given a graph $G=(V,E)$, the first homology group of $G$ is a free abelian group on a cycle basis of $G$. A cycle basis of $G$ is a minimal set of cyclic subgraphs of $G$ such that all cyclic subgraphs of $G$ can be expressed as unions or complements of elements in the cycle basis. For a more rigorous treatment on the construction of homology groups of topological spaces in general, we refer to Chapter 2 of  (Hatcher, 2002). of the given graph datasets. In combination with the novel positional embedding layer, we explicitly encode cyclic subgraphs in the dual-path message passing layer and incorporate channel-wise graph information in the graph information layer. These are combined with global attention across all Graph Transformer layers, a design of which was inspired from (Choi et al., 2023) and (Rampášek et al., 2023). As a result, the TIGT layer can concatenate hidden representations from the dual-path message passing layer, thereby combining information of the original graph dataset and its cyclic subgraphs, global attention layer, and graph information layer. This allows TIGT to preserve both topological information and graph-level information in each layer. Specifically, the dual-path message passing layer enables TIGT to overcome the limitations of positional encoding and structural encoding to increase expressive power when the number of layers increases. We justify the proposed model’s expressive power based on the theory of covering spaces. Furthermore, we perform experiments on synthetic datasets to distinguish isomorphism classes of graphs. We also evaluate the proposed model on benchmark datasets to demonstrate its state-of-the-art predictive performance.
@@ 2 Preliminary
Message passing neural networks. MPNNs have demonstrated proficiency in generating vector representations of graphs by exploiting local information based on node connectivity. This capability is shared with other types of GNNs, such as Graph Convolutional Network (GCN), Graph Attention Network (GAT) (Veličković et al., 2018), Graph Isomorphism Network (GIN) (Xu et al., 2019) and Residual Graph ConvNets (GatedGCN) (Bresson & Laurent, 2018). We use the abbreviation $\text{MPNN}^{l}$ to denote an MPNN that has a composition of $l$ neighborhood aggregating layers. Each $l$-th layer $H^{(l)}$ of the network constructs hidden node attributes of dimension $k_{l}$, denoted as $h_{v}^{(l)}$, using the following composition of functions:
@@
Let $M^{(L)}$ be the collection of all multisets of $k_{L}$-dimensional vectors with #$\#V(G)$ elements. Let
@@
Clique adjacency matrix The clique adjacency matrix, proposed by  (Choi et al., 2023), is a matrix that represents bases of cycles of a graph in a form analogous to the adjacency matrix, enabling its processing within GNNs. Extracting bases of cycles as in  (Paton, 1969), removing edges of the original graph not contained in the cycle bases, and substituting cyclic subgraphs in the cycle bases with cliques result in incorporating topological properties equivalent to first homological invariants (or cyclic substructures) of graphs. The set of cyclic subgraphs of $G$ which forms the basis of the cycle space (or the first homology group) of $G$ is defined as the cycle basis $\mathcal{B}_{G}$. The clique adjacency matrix, $A_{C}$, is the adjacency matrix of the union of #$\#\mathcal{B}_{G}$ complete subgraphs, each obtained from adding all possible edges among the set of nodes of each basis element $B\in\mathcal{B}_{G}$. Explicitly, the matrix $A_{C}:=\{a^{C}_{u,v}\}_{u,v\in V(G)}$ is given by
@@ 3 TOPOLOGY-INFORMED GRAPH TRANSFORMER(TIGT)
In this section, we introduce the overall TIGT architecture. The overall architecture of our model is illustrated in Figure 1. Suppose we are given the graph $G:=(V_{G},E_{G})$, where $V_{G}$ is the set of nodes and $E_{G}$ is the set of edges of $G$. The graph $G$ can be represented by four types of matrices which are used as inputs of TIGT: (1) a node feature matrix $X\in\mathbb{R}^{n\times k_{X}}$, (2) an adjacency matrix $A\in\mathbb{R}^{n\times n}$, (3) a clique adjacency matrix $A_{c}\in\mathbb{R}^{n\times n}$, and (4) an edge feature matrix $E\in\mathbb{R}^{n\times k_{E}}$. Note that $n$ is the number of nodes, $k_{X}$ is node feature dimension, and $k_{E}$ is edge feature dimension. The clique adjacency matrices are obtained using the same procedure outlined in previous research (Choi et al., 2023). For clarity and conciseness in our presentation, we leave the details pertaining to the normalization layer and the residual connection in Appendix C.2. We note that some of the mathematical notations used in explaining the model design and details of model structures conform to those shown in (Rampášek et al., 2023).
@@ 3.1 Topological positional embedding layer
where $X$ represents the node embedding tensor from the embedding layers, and [][\ \ ][ ] denotes the process of stacking two hidden representations. It’s important to note that MPNNs for the adjacency matrix and the clique adjacency matrix share weights. Then the node features are updated along with topological positional attributes, as shown below:
@@ 3.3 Mathematical background of TIGT
Positional encoding     As aforementioned, the method of imposing additional positional encodings to graph attributes can allow neural networks to distinguish cyclic structures, as shown in various types of Graph Transformers  (Ma et al., 2023; Rampášek et al., 2023; Ying et al., 2021). One drawback, however, is that these encodings may not be effective enough to represent classes of topologically non-isomorphic graphs as distinct vectors which are not similar to one another. We present a heuristic argument of the drawback mentioned above. Suppose we utilize a Transformer with finitely many layers to obtain vector representations $X^{*}_{G},X^{*}_{H}\in\mathbb{R}^{m}$ of two graphs $G$ and $H$ with node attribute matrices $X_{G},X_{H}\in\mathbb{R}^{n\times k}$ and positional encoding matrices $POS_{G},POS_{H}\in\mathbb{R}^{n\times k^{\prime}}$. Suppose further that all layers of the Transformer are comprised of compositions of Lipschitz continuous functions. This implies that the Transformer can be regarded as a Lipschitz continuous function from $\mathbb{R}^{n\times(k+k^{\prime})}$ to $\mathbb{R}^{m}$. Hence, for any $\epsilon>0$ such that $\|[X_{G}|POS_{G}]-[X_{H}|POS_{H}(v)]\|<\epsilon$, there exists a fixed constant $K>0$ such that $\|X^{*}_{G}-X^{*}_{H}\|<K\epsilon$. This suggests that if the node attributes and the positional encodings of non-isomorphic classes of graphs are similar to one another, say within $\epsilon$-error, then such Transformers will represent these graphs as similar vectors, say within $K\epsilon$-error. Hence, it is crucial to determine whether the given positional encodings effectively perturbs the node attributes to an extent that results in obtaining markedly different vector representations.
@@
Let $\mathcal{G}$ be any collections of graphs whose elements satisfy the following three conditions: (1) All graphs $G\in\mathcal{G}$ share the same number of nodes and edges, (2) any $G\in\mathcal{G}$ contains a cyclic subgraph with odd number of nodes, and (3) for any number $d\geq 1$, all the graphs $G\in\mathcal{G}$ have identical number of nodes whose degree is equal to $d$. Fix an integer $K$, and suppose the node indices for $G\in\mathcal{G}$ are ordered based on its increasing degrees. Let $\mathbf{P}$ be the RRWP positional encoding associated to $G$ defined as $\mathbf{P}_{i,j}:=[\textbf{I},\textbf{M},\textbf{M}^{2},\cdots,\textbf{M}^{K-1}]_{i,j}\in\mathbb{R}^{K}$, where $\textbf{M}:=\textbf{D}^{-1}\textbf{A}$ with A being the adjacency matrix of $G$, and D the diagonal matrix comprised of node degrees of $G$. Then there exists a unique vector $\pi\in\mathbb{R}^{n}$ independent of the choice of elements in $\mathcal{G}$ and a number $10<\gamma<$ 1 such that for any $10\leq l\leq K-$ 1, we have $\max_{(i,j)}\|\textbf{M}^{l}_{i,j}-\pi_{j}\|<\gamma^{l}$.
@@ 4 EXPERIMENTS
Models     To evaluate the discriminative power of TIGT, we compare a set of previous researches related to expressive power of GNNs on CSL dataset such as Graph Transformers (GraphGPS (Rampášek et al., 2023), GRIT (Ma et al., 2023)) and other message-passing neural networks (GCN (Kipf & Welling, 2017), GIN, Relational Pooling GIN(RP-GIN) (Murphy et al., 2019), Cy2C-GNNs (Choi et al., 2023)). We compare our approach on well-known benchmark datasets to test graph-level tasks with the latest SOTA techniques, widely adopted MPNNs models, and various Graph Transformer-based studies: GRIT (Ma et al., 2023), GraphGPS (Rampášek et al., 2023)), GCN (Kipf & Welling, 2017), GIN (Xu et al., 2019), its variant with edge-features (Hu et al., 2020), GAT (Veličković et al., 2018), GatedGCN (Bresson & Laurent, 2018), GatedGCN-LSPE (Dwivedi et al., 2022), PNA (Corso et al., 2020), Graphormer (Ying et al., 2021), K-Subgraph SAT (Chen et al., 2022), EGT (Hussain et al., 2022), SAN (Kreuzer et al., 2021), Graphormer-URPE (Luo et al., 2022), Graphormer-GD (Zhang et al., 2023), DGN (Beaini et al., 2021), GSN  (Bouritsas et al., 2021), CIN (Bodnar et al., 2021b), CRaW1  (Tönshoff et al., 2023), and GIN-AK+ (Zhao et al., 2022).
TIGT Setup     For hyperparameters of models on CSL datasets, we fixed the hidden dimension and batch size to 16, and other hyperparameters were configured similarly to the setting designed for the ZINC dataset. For a fair comparison of the other nine benchmark datasets, we ensured that both the hyperparameter settings closely matched those found in the GraphGPS (Rampášek et al., 2023) and GRIT (Ma et al., 2023) studies. The differences in the number of trainable parameters between TIGT and GraphGPS primarily arise from the additional components introduced to enrich topological information within the Graph Transformer layers. Further details on hyperparameters, such as the number of layers, hidden dimensions, and the specific type of MPNNs, are elaborated upon in the Appendix C.3.
@@
GCN 0.367$\pm$0.011 90.705$\pm$0.218 55.710$\pm$0.381 71.892$\pm$0.334 68.498$\pm$0.976
GIN 0.526$\pm$0.051 96.485$\pm$0.252 55.255$\pm$1.527 85.387$\pm$0.136 64.716$\pm$1.553
GAT 0.384$\pm$0.007 95.535$\pm$0.205 64.223$\pm$0.455 78.271$\pm$0.186 73.840$\pm$0.326
GatedGCN 0.282$\pm$0.015 97.340$\pm$0.143 67.312$\pm$0.311 85.568$\pm$0.088 73.840$\pm$0.326
GatedGCN+LSPE 0.090$\pm$0.001 N/A N/A N/A N/A
PNA 0.188$\pm$0.004 97.94$\pm$0.12 70.35$\pm$0.63 N/A N/A
DGN 0.168$\pm$0.003 N/A 72.838$\pm$0.417 86.680$\pm$0.034 N/A
GSN 0.101$\pm$0.010 N/A N/A N/A N/A
CIN 0.079$\pm$0.006 N/A N/A N/A N/A
CRaW1 0.085$\pm$0.004 97.944$\pm$0.050 69.013$\pm$0.259 N/A N/A
GIN-AK+ 0.080$\pm$0.001 N/A 72.19$\pm$0.13 86.850$\pm$0.057 N/A
SAN 0.139$\pm$0.006 N/A N/A 86.581$\pm$0.037 76.691$\pm$0.65
Graphormer 0.122$\pm$0.006 N/A N/A N/A N/A
K-Subgraph SAT 0.094$\pm$0.008 N/A N/A 86.848$\pm$0.037 77.856$\pm$0.104
EGT 0.108$\pm$0.009 98.173$\pm$0.087 68.702$\pm$0.409 86.821$\pm$0.020 79.232$\pm$0.348
Graphormer-GD 0.081$\pm$0.009 N/A N/A N/A N/A
GPS 0.070$\pm$0.004 98.051$\pm$0.126 72.298$\pm$0.356 86.685$\pm$0.059 78.016$\pm$0.180
GRIT 0.059$\pm$0.002 98.108$\pm$0.111 76.468$\pm$0.881 87.196$\pm$0.076 80.026$\pm$0.277
Cy2C-GNNs 0.102$\pm$0.002 97.772$\pm$0.001 64.285$\pm$0.005 86.048$\pm$0.005 64.932$\pm$0.003
TIGT 0.057$\pm$0.002 98.230$\pm$0.133 73.955$\pm$0.360 86.680$\pm$0.056 78.033$\pm$0.218

### v3

@@ 1 INTRODUCTION
Transformers have achieved remarkable success in domains such as Natural Language Processing (Vaswani et al., 2023) and Computer Vision (Dosovitskiy et al., 2021). Motivated by their prowess, researchers have applied them to the field of Graph Neural Networks (GNNs). They aimed to surmount the limitations of Message-Passing Neural Networks (MPNNs), a subset of GNNs, by attempting to address challenges such as over-smoothing (Oono and Suzuki, 2021), over-squashing (Alon and Yahav, 2021), and restricted expressive power (Xu et al., 2019; Morris et al., 2021). An exemplary application of the integration of Transformers into GNNs is the Graph Transformer. One way to achieve this integration is by applying multi-head attention mechanism of Transformers to each node of a given graph dataset. Such a technique treats the set of all nodes as being fully interconnected or connected by edges. This approach, however, often has limited capabilities in guaranteeing a strong inductive bias  [^1] For example, unless trained over large scale data, transformers may lack inherent inductive biases in modeling permutation and translational invariance, hierarchical structures, and local data structures  (Hahn, 2020; Dosovitskiy et al., 2021; Xu et al., 2021), hence making it prone to overfitting. To address this drawback, previous studies focused on devising new implementations to blend Graph Transformers with other deep learning techniques including MPNNs, thereby yielding promising empirical results  (Yang et al., 2021; Ying et al., 2021; Dwivedi and Bresson, 2021; Chen et al., 2022; Hussain et al., 2022; Zhang et al., 2023; Ma et al., 2023; rampášek2023recipe; Kong et al., 2023; Zhang et al., 2022).
@@
While Graph Transformers boast considerable advancements, enhancing them by strengthening their discriminative powers in distinguishing isomorphism classes of graphs still remains a challenge, an approach that can potentially boost their performance in predicting various properties of graph datasets. For instance, studies based on MPNN have enhanced node attributes by utilizing high-dimensional complexes, persistent homological techniques, and recurring subgraph structures (carrière2020perslay; Bodnar et al., 2021b; Bouritsas et al., 2021; Wijesinghe and Wang, 2021; Bevilacqua et al., 2022; Park et al., 2022; Horn et al., 2022; Choi et al., 2023; Li et al., 2023; Choi et al., 2024; Zhou et al., 2024). Other approaches addressing these limitations include incorporating positional encoding grounded in random walk strategies, Laplacian Positional Encoding (PE), node degree centrality, and shortest path distance. Furthermore, structure encoding based on substructure similarity has been introduced to amplify the inductive biases inherent in the Transformer.
@@
This paper introduces a Topology-Informed Graph Transformer (TIGT), which embeds topological inductive biases to augment the model’s expressive power and predictive efficacy. Prior to the Transformer layer, TIGT augments each node attribute with a topological positional embedding layer based on the differences of universal covers (or unfolding trees) obtained from the original graph and collections of unions of its cyclic subgraphs. These new topological biases allow TIGT to encapsulate the first homological invariants (or cyclic structures) [^2] Given a graph $G=(V,E)$, the first homology group of $G$ is a free abelian group on a cycle basis of $G$. A cycle basis of $G$ is a minimal set of cyclic subgraphs of $G$ such that all cyclic subgraphs of $G$ can be expressed as unions or complements of elements in the cycle basis. For a more rigorous treatment on the construction of homology groups of topological spaces in general, we refer to Chapter 2 of  (Hatcher, 2002). of the given graph datasets. In combination with the novel positional embedding layer, we explicitly encode cyclic subgraphs in the dual-path message passing layer and incorporate channel-wise graph information in the graph information layer. These are combined with global attention across all Graph Transformer layers, a design of which was inspired from (Choi et al., 2023) and (rampášek2023recipe). As a result, the TIGT layer can concatenate hidden representations from the dual-path message passing layer, thereby combining information of the original graph dataset and its cyclic subgraphs, global attention layer, and graph information layer. This allows TIGT to preserve both topological information and graph-level information in each layer. Specifically, the dual-path message passing layer enables TIGT to overcome the limitations of positional encoding and structural encoding to increase expressive power when the number of layers increases. We justify the proposed model’s expressive power based on the theory of covering spaces. Furthermore, we perform experiments on synthetic datasets to distinguish isomorphism classes of graphs. We also evaluate the proposed model on benchmark datasets to demonstrate its state-of-the-art predictive performance.
@@ 2 Preliminary
Message passing neural networks. MPNNs have demonstrated proficiency in generating vector representations of graphs by exploiting local information based on node connectivity. This capability is shared with other types of GNNs, such as Graph Convolutional Network (GCN), Graph Attention Network (GAT) (veličković2018graph), Graph Isomorphism Network (GIN) (Xu et al., 2019) and Residual Graph ConvNets (GatedGCN) (Bresson and Laurent, 2018). We use the abbreviation $\text{MPNN}^{l}$ to denote an MPNN that has a composition of $l$ neighborhood aggregating layers. Each $l$-th layer $H^{(l)}$ of the network constructs hidden node attributes of dimension $k_{l}$, denoted as $h_{v}^{(l)}$, using the following composition of functions:
@@
Let $M^{(L)}$ be the collection of all multisets of $k_{L}$-dimensional vectors with #V(G)\#V(G) elements. Let
@@
Clique adjacency matrix The clique adjacency matrix, proposed by  (Choi et al., 2023), is a matrix that represents bases of cycles of a graph in a form analogous to the adjacency matrix, enabling its processing within GNNs. Extracting bases of cycles as in  (Paton, 1969), removing edges of the original graph not contained in the cycle bases, and substituting cyclic subgraphs in the cycle bases with cliques result in incorporating topological properties equivalent to first homological invariants (or cyclic substructures) of graphs. The set of cyclic subgraphs of $G$ which forms the basis of the cycle space (or the first homology group) of $G$ is defined as the cycle basis $\mathcal{B}_{G}$. The clique adjacency matrix, $A_{C}$, is the adjacency matrix of the union of $\mathcal{B}_{G}$ complete subgraphs, each obtained from adding all possible edges among the set of nodes of each basis element $B\in\mathcal{B}_{G}$. Explicitly, the matrix $A_{C}:=\{a^{C}_{u,v}\}_{u,v\in V(G)}$ is given by
@@ 3 TOPOLOGY-INFORMED GRAPH TRANSFORMER(TIGT)
In this section, we introduce the overall TIGT architecture. The overall architecture of our model is illustrated in Figure 1. Suppose we are given the graph $G:=(V_{G},E_{G})$, where $V_{G}$ is the set of nodes and $E_{G}$ is the set of edges of $G$. The graph $G$ can be represented by four types of matrices which are used as inputs of TIGT: (1) a node feature matrix $X\in\mathbb{R}^{n\times k_{X}}$, (2) an adjacency matrix $A\in\mathbb{R}^{n\times n}$, (3) a clique adjacency matrix $A_{c}\in\mathbb{R}^{n\times n}$, and (4) an edge feature matrix $E\in\mathbb{R}^{n\times k_{E}}$. Note that $n$ is the number of nodes, $k_{X}$ is node feature dimension, and $k_{E}$ is edge feature dimension. The clique adjacency matrices are obtained using the same procedure outlined in previous research (Choi et al., 2023). For clarity and conciseness in our presentation, we leave the details pertaining to the normalization layer and the residual connection in Appendix C.2. We note that some of the mathematical notations used in explaining the model design and details of model structures conform to those shown in (rampášek2023recipe).
@@ 3.1 Topological positional embedding layer
where $X$ represents the node embedding tensor from the embedding layers, and [][\ \ ] denotes the process of stacking two hidden representations. It’s important to note that MPNNs for the adjacency matrix and the clique adjacency matrix share weights. Then the node features are updated along with topological positional attributes, as shown below:
@@ 3.3 Mathematical background of TIGT
Positional encoding     As aforementioned, the method of imposing additional positional encodings to graph attributes can allow neural networks to distinguish cyclic structures, as shown in various types of Graph Transformers  (Ma et al., 2023; rampášek2023recipe; Ying et al., 2021). One drawback, however, is that these encodings may not be effective enough to represent classes of topologically non-isomorphic graphs as distinct vectors which are not similar to one another. We present a heuristic argument of the drawback mentioned above. Suppose we utilize a Transformer with finitely many layers to obtain vector representations $X^{*}_{G},X^{*}_{H}\in\mathbb{R}^{m}$ of two graphs $G$ and $H$ with node attribute matrices $X_{G},X_{H}\in\mathbb{R}^{n\times k}$ and positional encoding matrices $POS_{G},POS_{H}\in\mathbb{R}^{n\times k^{\prime}}$. Suppose further that all layers of the Transformer are comprised of compositions of Lipschitz continuous functions. This implies that the Transformer can be regarded as a Lipschitz continuous function from $\mathbb{R}^{n\times(k+k^{\prime})}$ to $\mathbb{R}^{m}$. Hence, for any $\epsilon>0$ such that $\|[X_{G}|POS_{G}]-[X_{H}|POS_{H}(v)]\|<\epsilon$, there exists a fixed constant $K>0$ such that $\|X^{*}_{G}-X^{*}_{H}\|<K\epsilon$. This suggests that if the node attributes and the positional encodings of non-isomorphic classes of graphs are similar to one another, say within $\epsilon$-error, then such Transformers will represent these graphs as similar vectors, say within $K\epsilon$-error. Hence, it is crucial to determine whether the given positional encodings effectively perturbs the node attributes to an extent that results in obtaining markedly different vector representations.
@@
Let $\mathcal{G}$ be any collections of graphs whose elements satisfy the following three conditions: (1) All graphs $G\in\mathcal{G}$ share the same number of nodes and edges, (2) any $G\in\mathcal{G}$ contains a cyclic subgraph with odd number of nodes, and (3) for any number $d\geq 1$, all the graphs $G\in\mathcal{G}$ have identical number of nodes whose degree is equal to $d$. Fix an integer $K$, and suppose the node indices for $G\in\mathcal{G}$ are ordered based on its increasing degrees. Let $\mathbf{P}$ be the RRWP positional encoding associated to $G$ defined as $\mathbf{P}_{i,j}:=[\textbf{I},\textbf{M},\textbf{M}^{2},\cdots,\textbf{M}^{K-1}]_{i,j}\in\mathbb{R}^{K}$, where $\textbf{M}:=\textbf{D}^{-1}\textbf{A}$ with A being the adjacency matrix of $G$, and D the diagonal matrix comprised of node degrees of $G$. Then there exists a unique vector $\pi\in\mathbb{R}^{n}$ independent of the choice of elements in $\mathcal{G}$ and a number $0<\gamma<1$ such that for any $0\leq l\leq K-1$, we have $\max_{(i,j)}\|\textbf{M}^{l}_{i,j}-\pi_{j}\|<\gamma^{l}$.
@@ 4 EXPERIMENTS
Models     To evaluate the discriminative power of TIGT, we compare a set of previous researches related to expressive power of GNNs on CSL dataset such as Graph Transformers (GraphGPS (rampášek2023recipe), GRIT (Ma et al., 2023)) and other message-passing neural networks (GCN (Kipf and Welling, 2017), GIN, Relational Pooling GIN(RP-GIN) (Murphy et al., 2019), Cy2C-GNNs (Choi et al., 2023)). We compare our approach on well-known benchmark datasets to test graph-level tasks with the latest SOTA techniques, widely adopted MPNNs models, and various Graph Transformer-based studies: GRIT (Ma et al., 2023), GraphGPS (rampášek2023recipe)), GCN (Kipf and Welling, 2017), GIN (Xu et al., 2019), its variant with edge-features (Hu et al., 2020), GAT (veličković2018graph), GatedGCN (Bresson and Laurent, 2018), GatedGCN-LSPE (Dwivedi et al., 2022), PNA (Corso et al., 2020), Graphormer (Ying et al., 2021), K-Subgraph SAT (Chen et al., 2022), EGT (Hussain et al., 2022), SAN (Kreuzer et al., 2021), Graphormer-URPE (Luo et al., 2022), Graphormer-GD (Zhang et al., 2023), DGN (Beaini et al., 2021), GSN  (Bouritsas et al., 2021), CIN (Bodnar et al., 2021b), CRaW1  (tönshoff2023walking), and GIN-AK+ (Zhao et al., 2022).
TIGT Setup     For hyperparameters of models on CSL datasets, we fixed the hidden dimension and batch size to 16, and other hyperparameters were configured similarly to the setting designed for the ZINC dataset. For a fair comparison of the other nine benchmark datasets, we ensured that both the hyperparameter settings closely matched those found in the GraphGPS (rampášek2023recipe) and GRIT (Ma et al., 2023) studies. The differences in the number of trainable parameters between TIGT and GraphGPS primarily arise from the additional components introduced to enrich topological information within the Graph Transformer layers. Further details on hyperparameters, such as the number of layers, hidden dimensions, and the specific type of MPNNs, are elaborated upon in the Appendix C.3.
@@
[…]
GSN $1±\pm0.0$10 N/A N/A N/A N/A
[…]